Find the slope of any line parallel to the line through points (15, 1) and (4, 2).
Answer by checkley71(8403)

(2-1)/(4-15)
1/-11
-1/11 IS THE SLOPE
Y=mX+b
1=(-1/11)15+b
1+15/11=b
b=(11+15)/11
b=26/11
THUS THE EQUATION OF THIS LINE IS
Y=-X/11+26/11